OpenStudy (anonymous):

The distance from Tampa to Lake Wales is 55 miles. The distance from Tampa to Dunnellon is about 45% longer. To the nearest mile, what is the distance between Tampa and Dunnellon?

OpenStudy (anonymous):

45% in decimal form is 0.45. The distance from Tampa to Dunnellon would be expressed as\[d = 55 + 55*0.45\] or\[d = 55(1.45)\]